# ReGUIDE: Data Efficient GUI Grounding via Spatial Reasoning and Search

### v1

@@ 2 Related Work
Test-time scaling. Recent works explored that scaling test-time computation, such as best-of-N sampling, can be even better than scaling train-time computation for performance [29]. Specifically, test-time scaling strategies improve LLM performance by generating numerous candidate outputs and selecting the best. To enhance decision-making, external verifiers are often employed to evaluate and refine these outputs [14]. In the localization task, recent work framed the localization as a search problem and suggested a test-time scalable strategy [31].
@@ 4 Experiments
Before answering each question, we outline the experimental protocol (more details in Appendix).
@@
Training setup. For the main experiment, we train ReGUIDE on Qwen-2.5-VL 3B/7B. To demonstrate ReGUIDE’s data efficiency, we utilize only a 20k subset of UGround [10] dataset constituting approximately 0.2% of its full set. Learning to explain web images via reasoning, we use the GRPO [8] algorithm with two rewards, accuracy reward and output-format compliance reward, and $\lambda=0.1$ for formatting reward weight. For learning to predict consistent coordinates under transformations, where the local view is a random crop of up to 30% of the original area while preserving aspect ratio. More setups in Appendix.
@@ 4.1 Main Results
As shown in Table 1, we present the GUI grounding performance of ReGUIDE compared to other baselines. We first compare models trained with the same base architecture (Qwen-2.5-VL) and dataset (a 20k subset of UGround). Our method, ReGUIDE, significantly outperforms supervised fine-tuned (SFT) baselines across all evaluated scenarios. For instance, ReGUIDE achieves an accuracy of 87.3% on ScreenSpot with the 3B model. Similar improvements are observed for the more challenging ScreenSpot-Pro benchmark, where ReGUIDE boosts performance from 23.9% to 44.3%. Also in ScreenSpot-v2, ReGUIDE outperforms baselines with a meaningful margin.
@@ 4.2 Additional Analysis and Ablation
Analysis of training components. We perform an analysis on each component of ReGUIDE, specifically spatial reasoning (Reasoning), learn consistency under transformation (Consistency), and test-time searching strategy (Searching). As shown in Table 4, each component plays an important role, leading to gradual and significant improvements when applied sequentially. Additional comparisons of alternative RL policy-optimization algorithms are provided in Appendix.
@@ 4.3 Agentic Settings Experiments
To verify whether better grounding transfers to real tasks, we plug each grounding model with ReGUIDE and use a high-level planner (GPT-4o). Here, we report task-success rate (%) on five offline suites: Multimodal-Mind2Web-cross domain, -cross task, and -cross website [9], AndroidControl-Low and -High [18]. Each benchmark has a different domain of agent settings (i.e., desktop web navigation and mobile-app control). For every episode, the planner emits a natural-language action plan; the grounding module converts each step into pixel coordinates, which the environment executes. Success is recorded when the episode reaches the goal state. All prompts, settings, and frameworks are followed by UGround [10].

### v2

@@ 2 Related Work
Test-time scaling. Recent works explored that scaling test-time computation, such as best-of-N sampling, can be even better than scaling train-time computation for performance [32]. Specifically, test-time scaling strategies improve LLM performance by generating numerous candidate outputs and selecting the best [32, 19, 15]. To enhance decision-making, external verifiers are often employed to evaluate and refine these outputs [15]. In the localization task, recent work framed the localization as a search problem and suggested a test-time scalable strategy [34].
@@ 4 Experiments
Before answering each question, we outline the experimental protocol (more details in Appendix A).
@@
Training setup. For the main experiment, we train ReGUIDE on Qwen-2.5-VL 3B/7B. To demonstrate ReGUIDE’s data efficiency, we utilize only a 20k subset of UGround [11] dataset constituting approximately 0.2% of its full set. Learning to explain web images via reasoning, we use the GRPO [9] algorithm with two rewards, accuracy reward and output-format compliance reward, and $\lambda=0.1$ for formatting reward weight. For learning to predict consistent coordinates under transformations, where the local view is a random crop of up to 30% of the original area while preserving aspect ratio. More setups in Appendix A.3.
@@ 4.1 Main Results
As shown in Table 1, we present the GUI grounding performance of ReGUIDE compared to other baselines. We first compare models trained with the same base architecture (Qwen-2.5-VL) and dataset (a 20k subset of UGround). Our proposed method, ReGUIDE, significantly outperforms supervised fine-tuning (SFT) baselines across all evaluated scenarios. For instance, ReGUIDE achieves an accuracy of 87.3% on ScreenSpot with the 3B model. Similar improvements are observed for the more challenging ScreenSpot-Pro benchmark, where ReGUIDE boosts performance from 23.9% to 44.3%. Also in ScreenSpot-v2 and dataset, ReGUIDE outperforms baseline and SFT model with a meaningful margin.
@@ 4.2 Additional Analysis and Ablation
Analysis of training components. We perform an analysis on each component of ReGUIDE, specifically spatial reasoning (Reasoning), learn consistency under transformation (Consistency), and test-time searching strategy (Searching). As shown in Table 4, each component plays an important role, leading to gradual and significant improvements when applied sequentially. Additional comparisons of alternative RL policy-optimization algorithms are provided in Appendix D.
@@ 4.3 Agentic Settings Experiments
To verify whether better grounding transfers to real tasks, we plug each grounding model with ReGUIDE and use a high-level planner (GPT-4o). Here, we report task-success rate (%) on five offline suites: Multimodal-Mind2Web-cross domain, -cross task, and -cross website [10], AndroidControl-Low and -High [21]. Each benchmark has a different domain of agent settings (i.e., desktop web navigation and mobile-app control). For every episode, the planner emits a natural-language action plan; the grounding module converts each step into pixel coordinates, which the environment executes. Success is recorded when the episode reaches the goal state. All prompts, settings, and frameworks are followed by UGround [11]. The detailed settings are described in Appendix A.
@@ 5 Discussion and Conclusion
Appendix A Experimental Details
In this section, we describe the experimental details of Section 4, including ReGUIDE and baselines.
A.1 Dataset Details
In this section, we describe the dataset we used in training and evaluation.
  • UGround Uground is a Graphical User Interface (GUI) grounding dataset that consists of 10M triplets of GUI image, element coordinates, and description of the element, which was synthesized by open MLLM, Llava-Next-13 B. Uground [11] models are trained with their own synthesized web GUI dataset, which consists of 10 M image-instruction pairs. For training ReGUIDE, we randomly select a 20K image-instruction set from UGround [11] dataset, which is approximately 0.2% of the original dataset.
  • ScreenSpot ScreenSpot [7] is an evaluation benchmark dataset which consists of 1,272 (image, description, bounding box) triplets. There are various GUI domains in multiple device environments (Mobile, Desktop, Web) and UI element types (Text, Icon).
  • ScreenSpot-v2 Wu et al. [35] refined the original ScreenSpot dataset by resolving ambiguous descriptions and aligning annotations more precisely with visual elements. We use both versions for a more accurate grounding evaluation.
  • ScreenSpot-pro ScreenSpot-Pro [20] is a benchmark designed to evaluate GUI grounding models in professional, high-resolution environments. It spans 23 applications across five professional categories and three operating systems, highlighting the challenges models face when interacting with complex software..
  • Android Control Android Control [21] is an offline agent benchmark comprising 15000 distinct user tasks sampled from 833 different Android applications. Each task contains a sequence of screenshots, detailed action records, and accessibility trees derived from human demonstrations (the “golden trajectory”). Each action is also annotated with a specific instruction (e.g., “set the hours to 6”), allowing both high-level and low-level task settings. In our experiments, we followed Gou et al. [11] and evaluated models on 500 randomly sampled steps from the test split.
  • Multimodal-Mind2Web The Multimodal-Mind2Web [10] dataset contains screenshots with a large vertical dimension (e.g., 1280 x 10000 pixels). The test split includes 1,013 realistic user tasks collected from over 100 different websites, each accompanied by a high-level instruction and a sequence of action steps. To make the data manageable we followed [11], and divide the image vertically with certain amount of duplication. During the agent evaluation, when an agent cannot find an actionable element or explicitly chooses to scroll, the next block is processed, simulating user scrolling.
A.2 Model Details
  • Qwen-2.5-VL We use Qwen-2.5-VL [2] is a multimodal large language model (mllm) which inputs text and images and outputs text. The model is equipped with vision-language fusion modules and trained on diverse visual instruction datasets. We trained ReGUIDE from Qwen-2.5-VL-3B and 7B models.
  • UGround UGround [11] is the model trained from Qwen-2-VL [2], there are two versions UGround-2B and UGround-7B. They are trained with a synthesized dataset that has 10M images, instructions, and target coordinates triplets. This amount of dataset size is about 500 times bigger than the ReGUIDE trained. For the fair comparison, we trained UGround on Qwen-2.5-VL [2] with the same subset used to train ReGUIDE. As shown in Table 1, ReGUIDE excessively outperforms UGround for both.
  • Proprietary Models Endpoint
    • gpt-4o-2024-05-13
    • claude-3-7-sonnet-20250219
A.3 Training and Evaluation Details
  • Training framework For the training framework, we utilize VERL for reinforcement learning and TRL for fine-tuning.
    • TRL (https://github.com/huggingface/trl)
    • VERL (https://github.com/volcengine/verl)
  • Training hyperparameters. We describe our training hyperparameters, which we used for training, especially for learning to explaining GUI images via reasoning and learning to predict consistent coordinates under transformations.
    Table 7: Hyperparameters for ReGUIDE on grpo training
    Hyperparmeter Value
    Optimizer Adam
    Algorithm GRPO
    Learning rate 1e-6
    Training data size 20k
    Batch size 128
    Generation per Sample 8
    kl_coef 0.01
    Table 8: Hyperparameters for ReGUIDE on training consistent coordinated under transformations
    Hyperparmeter Value
    Optimizer Adam
    Learning rate 1e-6
    Training data size 20k
    Epoch 1
    Batch size 64
    minimum crop ratio 0.3
A.4 Compute Resource
For the main development, we mainly use Intel(R) Xeon(R) Gold 6338 CPU @ 2.00GHz and four A100 80GB GPUs. For training ReGUIDE, it took around 20 hours.
Appendix B Additional Experimental Results
B.1 Comparison with additional Baselines
Table 9: Accuracy (%) for ReGUIDE (Ours) and other baselines, including UI-TARS [28], UI-R1 [23], GUI-R1 [36], and ReGUIDE (Ours). We evaluate on three web-grounding benchmarks: ScreenSpot [7], ScreenSpot-v2 [35], and ScreenSpot-Pro [20].
Methods ScreenSpot [7] ScreenSpot-v2 [35] ScreenSpot-Pro [20] Average
UI-TARS-2B [28] 82.3 84.7 27.7 64.9
UI-TARS-7B [28] 89.5 91.6 35.7 72.3
UI-R1-E-3B [23] 89.2 89.5 33.5 70.7
GUI-R1-7B [36] - - 31.3 -
ReGUIDE-3B (Ours) 87.3 89.7 44.3 73.8
ReGUIDE-7B (Ours) 90.1 91.5 44.4 75.3
We compare the grounding accuracy of ReGUIDE with other additional baselines, which are closed-source (UI-TARS [28]) and the concurrently proposed reinforcement learning methods for grounding UI-R1 [23] and GUI-R1 [36]. As shown in Table 9 ReGUIDE achieves the highest average combined accuracy (75.3%) across ScreenSpot [7], ScreenSpot-V2 [35], and ScreenSpot-Pro [20].
Crucially, the compact ReGUIDE-3B already surpasses the best 7B-parameter base model (UI-TARS-7B) by +1.5 pp in average accuracy (72.3 % $\rightarrow$ 73.8 %), demonstrating that our training and search strategy substantially improves the GUI grounding performance. Moreover, both reinforcement fine-tuned baselines (i.e., UI-R1-E-3B, GUI-R1-7B) perform worse than ReGUIDE, underscoring the effectiveness of our further training for consistency under transformation and search algorithm. The full-sized ReGUIDE-7B further extends this lead to +1.5 pp, establishing a best among the given baselines on the combined benchmarks.
B.2 Effectiveness of Kernel Density Estimation as majority voting strategy
Table 10: Comparison of voting algorithm with grounding accuracy (%) on ScreenSpot [7] and ScreenSpot-Pro [20]
Methods ScreenSpot [7] ScreenSpot-Pro [20]
Center 80.3 20.2
Medoid 84.6 29.4
KDE (Ours) 87.3 44.3
To verify the most effective method for aggregating diverse pixel-wise coordinate predictions, we conducted an ablation study comparing our Kernel Density Estimation (KDE) based strategy against "Center" and "Medoid". "Center" is simply getting the mean point of the predicted coordinates. "Medoid" selects an actual predicted point that minimizes the sum of distances to all other predictions. As shown in Table 10, KDE shows notable results compared with other voting strategies.
The performance gap can be explained by the statistical properties of each strategy. First, the "Center" is vulnerable to outliers, a few erroneous predictions can shift the center far from the true target. Also, "Medoid" is robust to outliers but ignores local density, so it may choose a point in a sparse region when the prediction distribution is multi-modal. However, KDE can overcome these limitations. KDE estimating the underlying likelihood surface, naturally capturing local density peaks while down-weighting isolated outliers. This behaviour aligns with our empirical observation that model confidence around the true GUI element is approximately Gaussian (Figure 2), making KDE a principled and effective spatial voting rule.
B.3 Effect of consistency finetuning after reinforcement stage
Table 11: Effect of consistency-under-transformation finetuning on ReGUIDE-3B, evaluated on the ScreenSpot [7] benchmark.
Method Augmentation usage Accuracy (%)
RL w/o Aug. 83.3
RL w/ Aug. In RL stage 83.2
ReGUIDE (Ours). fine-tune 84.9
We assess the benefit of our consistency-under-transformation finetuning on the ReGUIDE-3B model using the ScreenSpot benchmark. As shown in Table 11, training GRPO on the original screenshots only (RL w/o Aug.) achieves 83.3 % accuracy, while mixing the same augmented views directly into the RL phase (RL w/ Aug) yields no improvement (83.2 %). In contrast, our two-stage pipeline—first training GRPO without augmentation and then applying a separate consistency finetune on the augmented views (ReGUIDE)—raises accuracy to 84.9 %, a gain of +1.6 pp over the vanilla RL model and +1.7 pp over RL w/ Aug. This result confirms that the performance boost arises from explicitly learning to align predictions across transformed views, not from data augmentation alone.
B.4 Effect of consistency finetuning on transformer attention
(a) GRPO trained
(b) ReGUIDE
Figure 4: Attention map between text tokens, X-axis is key tokens, Y-axis is query tokens. In the image, I denotes instruction, R denotes reasoning rationales, and A denotes Answer.
To assess how the consistency-under-transformation stage alters internal representations, we visualised the average attention logits of the last transformer layer in ReGUIDE-3B. We first removed turn-delimiter and <eos> tokens whose weights dominate the map, and then averaged across all heads. As shown in Figure 4, the GRPO-only model weights most of its attention onto the reasoning span, indicating limited contextual grounding. In contrast, the full ReGUIDE model, after consistency finetuning, distributes attention more evenly across reasoning (i.e., text tokens between <think> and </think>) and other text tokens, suggesting richer spatial awareness and reduced attention collapse.
B.5 Dataset Scalable Experiment
Figure 5: Grounding accuracy (%) on ScreenSpot [7] as the UGround [11] training set grows from 3.2 K to 20 K samples. We trained from Qwen-2.5-VL-3B.
To measure how our reinforcement-learning stage benefits from more data, we varied the amount of UGround [11] training samples from 3.2K up to 20K in 3.2K increments and evaluated grounding accuracy (%) on the ScreenSpot [7] benchmark (Figure 5). Accuracy rises monotonically with data size, and the full 20K subset yields the best performance. These results indicate that the ReGUIDE training pipeline scales effectively with additional data and has not yet saturated at the 20K level.
Appendix C Illustration of Searching Strategy.
Figure 6: Overview of ReGUIDE’s searching strategy.. Starting from the full screenshot, the model samples $N$ coordinate predictions, votes using kernel density estimation, and recenters a crop on the density peak. One searching pass is then run inside the crop, and the point with the highest KDE density is returned as the final prediction.
Appendix D Alternative RL policy optimization algorithms
Table 12: Comparison of RL algorithms. We report accuracy (%) and average output length (Output Len.) on ScreenSpot Benchmark [7].
Methods Accuracy (%) Output Len.
REINFORCE++ [16] 64.8 32.3
RLOO [1] 77.8 95.7
GRPO [9] 83.3 80.8
While we primarily adopt GRPO [9] for the learning to visual reasoning, alternative RL algorithms, such as RLOO [1], REINFORCE++ [16], could also be considered. To verify the performance difference between RL poly optimization algorithms, we compared grounding accuracy (%) on ScreenSpot. As shown in the Table 12, GRPO achieves the highest accuracy (83.3%), followed by RLOO (77.8%), while REINFORCE++ significantly underperforms (64.8%). Interestingly, we observed that the REINFORCE++ trained model loses its reasoning ability (i.e., shortest output length), eventually failing to generate meaningful reasoning paths.
Appendix E Learning Curve
We report the learning curve of ReGUIDE during training GRPO. The base model of the training curve is Qwen-2.5-VL-3B. We can see that the reward is kept increasing during the training. Also, the mean response length has decreased during the training.
(a) Mean reward score per step
(b) Mean response length per step
Figure 7: Learning curve during training GRPO of ReGUIDE.
Appendix F Prompts
We describe our generation prompt for ReGUIDE (in Table 13 and UGround [11] (in Table 14. We used the ReGUIDE prompt for training and evaluation ReGUIDE, and the UGround prompt for supervised fine-tuning the Qwen-2.5-VL for comparison and evaluating the UGround model.
Table 13: Prompt of ReGUIDE
Element: Target element instruction. size_x: Width of the given image. size_y: Height of the given image.
What is the coordinate of [{Element}] in the image? The size of image is ({size_x},{size_y}). Output the thinking process in <think> </think> and final answer (coordinate (x,y)) in <answer> </answer> tags.
Table 14: Prompt of UGround
Element: Target element instruction.
Your task is to help the user identify the precise coordinates (x, y) of a specific area/element/object on the screen based on a description. - Your response should aim to point to the center or a representative point within the described area/element/object as accurately as possible. - If the description is unclear or ambiguous, infer the most relevant area or element based on its likely context or purpose. - Your answer should be a single string (x, y) corresponding to the point of the interest. Description: {Element} Answer:
Appendix G Generation Examples
Table 15: Examples of ReGUIDE’s generated reasoning and predicted coordinate.
Target Prompt: add description. Response: <think>To find the coordinate of [add description], I need to locate the element labeled Ädd descriptionïn the image. The coordinates are given as (x, y) pairs, and I need to identify the position of this element.</think><answer>(300,2221) </answer>
Target Prompt: check weather on tuesday. Response: <think> To determine the coordinate of the n̈lpf̈older, I need to locate it in the image and identify its position relative to the other folders. The n̈lpf̈older is the second folder in the top row, as indicated by its position in the grid layout. </think> <answer>(584,760)</answer>
Target Prompt: sign up for newsletter. Response: <think> The task is to find the coordinate of the S̈ign Up for Newsletterc̈heckbox in the image. The checkbox is located in the P̈ersonal Informations̈ection, and it is a small square with a label next to it. The coordinates are given in a specific size of the image, and we need to identify the position of the checkbox within the image. </think> <answer>(367,85)</answer>
Target Prompt: start mapping. Response: <think> The S̈tart Mappingb̈utton is typically used to begin the process of adding or editing data on the map. In the image, it is located in the lower left section, which is a common placement for such buttons on web interfaces. </think> <answer>(558,805)</answer>
Appendix H Societal Impact
This paper presents ReGUIDE, a method that significantly improves GUI-grounding accuracy. We expect that our approach will enhance digital accessibility and productivity by enabling agents to locate and interact with on-screen elements more reliably. However, the same capability could be misused for malicious purposes, such as automated hacking or spreading misinformation, issues that current agent models already face. Mitigation measures include responsible release practices, access controls, and detection tools to discourage adversarial deployments.